Please computer typed not hand writing! Thanks
Answer #1

Part A Given mean x-500, sd(ơ-12, n-50 С 0.95 a 1-0.95-0.05 z-Zoes -1.96 (using z value table) 95% confidence interval Margin

Part B Given meanx-120, sd(s) -15, n-51 C-0.95 a 1-0.95-0.05 z-zoos-1.96 (using z value table) 95% confidence interval Margin of error (116823 Lowe limitx E 120-4.116823 115.8832 Upper limit E-120+ 4.116823 - 124.1168 95% confidence interval is (115.8832. 124.1168) V51 Using excel the answer is 120 (NORMSINV (0.05)*15/sqrt (51) (115.8833,124.1167)(-0.99 α= 1-0.99-0.01 z-Zoo1-2.58 (using z value table) 99% confidence interval Margin of error ( 258419084 Lowe limit-x-E = 120-5.419084 = 1 14.5 809 Upper limit-x + E = 120 + 5.4 19084-125.4191 ง่าเ 99% confidence interval is (1 14.5809. Î25.4 191) Using excel the answer is -120 (NORMSINV (0.01)#15/sqrt(SI)):(114.5897 , 125.4103)